# Recent progresses in strange quark stars

### v1

@@ Abstract
According to the hypothesis that strange quark matter may be the true ground state of matter at extremely high densities, strange quark stars should be stable and could exist in the Universe. It is possible that pulsars may actually be strange stars, but not neutron stars. Here we present a short review on recent progresses in the field of strange quark stars. Three popular phenomenological models widely used to describe strange quark matter are introduced, with special attention being paid on the corresponding equation of state in each model. Combining the equation of state with the Tolman-Oppenheimer-Volkov equations, the inner structure and mass-radius relation can be obtained for the whole sequence of strange stars. Strong gravitational wave emissions may be generated by strange stars through various mechanisms, which may help identify strange stars via observations. Especially, close-in strange quark planets with respect to their hosts may provide a unique test for the existence of strange quark objects. Fierce electromagnetic bursts could also be generated by strange stars. The energy may come from the phase transition of neutron stars to strange stars, or from the merger of binary strange stars. The collapse of the strange star crust can also release a huge amount of energy. It is shown that strange quark stars may be involved in short gamma-ray bursts and fast radio bursts.
@@ 3 Introduction
Interestingly, according to the strange quark matter hypothesis, strange quark dwarfs and even strange quark planets could also exist. They may be produced due to the contamination of white dwarfs/planets by strange nuggets in the Universe. While the stability of these low-mass strange objects is still highly debated (Glendenning et al., 1995; Vartanyan et al., 2012, 2014; Alford et al., 2017; Di Clemente et al., 2023; Gonçalves et al., 2023), they could provide valuable opportunities for identifying strange stars due to their significant difference from normal matter white dwarfs and planets (Huang and Yu, 2017; Kuerban et al., 2020; Wang et al., 2021; Kurban et al., 2022b).
The study of strange stars is an active and rapidly developing field. In the past few decades, various observational aspects of strange stars have been studied. In this review, we are going to present a brief description on some recent progresses concerning strange stars. The structure of our paper is organized as follows. The composition and internal structure of strange stars are introduced in Section 4. Especially, several popular phenomenological models describing the strong interaction among quarks are presented. In Section 5, gravitational wave (GW) emissions from various strange quark objects are discussed. Possible connection between some violent electromagnetic bursts (e.g. gamma-ray bursts and fast radio bursts) and strange stars are introduced in Section 6. Finally, Section 7 presents the conclusions and some further discussion.
@@ 4 Internal structure of strange stars
While thousands of pulsars have been observed, the internal composition and structure of them are still controversial. This enigma is closely connected with the interaction and properties of matter under extreme densities and temperatures, thus is an important issue involving fundamental physics. Theoretically, pulsars could be neutron stars or quark stars. In the quark star case, they could either be three-flavor (u, d, s) strange stars, or could even be two-flavor (u, d) quark stars. The existence of the so called hybrid stars is also suggested, which include a quark matter core encompassed by normal nuclear matter (Ivanenko and Kurdgelaidze, 1965; Annala et al., 2020; Menezes, 2021).
@@ 4.2 NJL model
The Nambu–Jona-Lasinio (NJL) model is initially proposed by Nambu and Jona-Lasinio (1961a, b) to describe the interaction between nucleons. It was extended by Eguchi and Sugawara (1974) to include up and down quarks. Kikkawa (1976) further developed the theory to encompass three flavors of quarks, including up, down, and strange quarks. In this review, we utilize the three-flavor NJL model to describe quark matter inside strange stars. The model exhibits spontaneous breaking of chiral symmetry, which is essential for understanding the large nucleon mass and the dynamic generation of fermion masses. Additionally, the model is noteworthy for its solvability, allowing for simple analytical results obtained in certain limiting cases.
@@ 4.3 Quasi-particle model
In view of quark confinement, the energy density in the vacuum is lower than that of free quarks. Consequently, the vacuum pressure at zero chemical potential ( ) is negative. Using the MIT bag model for reference, we can set . However, similar to the MIT bag model and NJL model, the quasi-particle model also cannot accurately describe the characteristics of asymptotic freedom.
4.4 Properties of strange stars
EOS can be derived from the above three popular quark matter models. Using the EOS, we can go further to determine the mass and radius of strange stars by solving the Tolman–Oppenheimer–Volkov (TOV) equations (Tolman, 1939; Oppenheimer and Volkoff, 1939), i.e.
@@ 4.4 Properties of strange stars
where is the gravitational constant, and represent the radius and mass, respectively. Typical mass-radius curves of bare strange stars derived by using the MIT bag model are shown in the left panel of Figure 1 (Deb et al., 2017). On the other hand, the mass-radius curves of crusted strange stars are shown in the right panel of Figure 1 (Huang and Lu, 1997). This mass-radius relationship help us understand the maximum and minimum limits of mass, the upper limit of radius, the density distribution and even the stability condition of strange stars. We can also compare the theoretically predicted parameters with observational data to test the model and gain insights into the nature of matter under extreme conditions.
The tidal deformability, which features the distortion of an object caused by external gravitational forces, can be measured through gravitational wave observations. It is quantified by the tidal Love number as (Thorne, 1998)
where is the dimensionless tidal deformation parameter. The detailed expression for has been given by Thorne (1998) and Flanagan and Hinderer (2008). The sound velocity () is another important parameter that reflects the stiffness of the EOS. It can be calculated as
4.5 Stellar stability
@@
The TOV equations determine the hydrostatic equilibrium of the stellar structure, while whether the resulting configuration is stable or not should be further investigated. In most cases, the stellar stability against radial oscillations is investigated by applying the Bardeen-Thorne-Meltzer (BTM) criterion (Bardeen et al., 1966). The BTM criterion is usually expressed as follows: when moving toward the direction of increasing central pressure along the mass-radius curve, at each extremum, one previously stable radial mode becomes unstable if the curve bends counterclockwise, while one previously unstable radial mode becomes stable if the curve bends clockwise. For bare strange quark objects which include the whole bare strange planet-bare strange star series, the conclusion on the stability is quite clear. All configurations before the mass-radius curve reaches its maximum are stable, with stars of higher central pressure unstable. It means that all the bare strange planets and bare strange dwarfs are stable.

### v2

@@ Abstract
According to the hypothesis that strange quark matter may be the true ground state of matter at extremely high densities, strange quark stars should be stable and could exist in the Universe. It is possible that pulsars may actually be strange stars, but not neutron stars. Here we present a short review on recent progresses in the field of strange quark stars. First, three popular phenomenological models widely used to describe strange quark matter are introduced, with special attention being paid on the corresponding equation of state in each model. Combining the equation of state with the Tolman-Oppenheimer-Volkov equations, the inner structure and mass-radius relation can be obtained for the whole sequence of strange stars. Tidal deformability and oscillations (both radial and non-radial oscillations), which are sensitive to the composition and the equations of state, are then described. Hybrid stars as a special kind of quark stars are discussed. Several other interesting aspects of strange stars are also included. For example, strong gravitational wave emissions may be generated by strange stars through various mechanisms, which may help identify strange stars via observations. Especially, close-in strange quark planets with respect to their hosts may provide a unique test for the existence of strange quark objects. Fierce electromagnetic bursts could also be generated by strange stars. The energy may come from the phase transition of neutron stars to strange stars, or from the merger of binary strange stars. The collapse of the strange star crust can also release a huge amount of energy. It is shown that strange quark stars may be involved in short gamma-ray bursts and fast radio bursts.
@@ 3 Introduction
Interestingly, according to the strange quark matter hypothesis, strange quark dwarfs and even strange quark planets could also exist. They may be produced due to the contamination of white dwarfs/planets by strange nuggets in the Universe. While the stability of these low-mass strange objects is still highly debated (Glendenning et al., 1995; Fraga et al., 2001; Vartanyan et al., 2012, 2014; Alford et al., 2017; Di Clemente et al., 2023; Gonçalves et al., 2023), they could provide valuable opportunities for identifying strange stars due to their significant difference from normal matter white dwarfs and planets (Huang and Yu, 2017; Kuerban et al., 2020; Wang et al., 2021; Kurban et al., 2022b).
The study of strange stars is an active and rapidly developing field. In the past few decades, various observational aspects of strange stars have been studied. In this review, we are going to present a brief description on some recent progresses concerning strange stars. Hybrid stars, in which quark matter and hadronic matter may coexist, are also discussed. The structure of our paper is organized as follows. The properties and internal structure of strange stars are introduced in Section 4. Especially, several popular phenomenological models describing the strong interaction among quarks are presented. Tidal deformability and oscillations, which are closely related to the internal structure, are also introduced. As a special kind of quark stars, hybrid stars are discussed in Section 5, paying special attention on the transition between different phases. In Section 6, gravitational wave (GW) emissions from various strange quark objects are discussed. Possible connection between some violent electromagnetic bursts (e.g. gamma-ray bursts and fast radio bursts) and strange stars are introduced in Section 7. Finally, Section 8 presents the conclusions and some further discussion.
@@ 4 Internal structure of strange stars
While thousands of pulsars have been observed, the internal composition and structure of them are still controversial. This enigma is closely connected with the interaction and properties of matter under extreme densities and temperatures, thus is an important issue involving fundamental physics. Theoretically, pulsars could be neutron stars or quark stars. In the quark star case, they could either be three-flavor (u, d, s) strange stars, or could even be two-flavor (u, d) quark stars. The existence of the so called hybrid stars is also suggested, which usually include a quark matter core encompassed by normal nuclear matter (Ivanenko and Kurdgelaidze, 1965; Annala et al., 2020; Menezes, 2021).
@@ 4.2 NJL model
The Nambu-Jona-Lasinio (NJL) model is initially proposed by Nambu and Jona-Lasinio (1961a, b) to describe the interaction between nucleons. It was extended by Eguchi and Sugawara (1974) to include up and down quarks. Kikkawa (1976) further developed the theory to encompass three flavors of quarks, including up, down, and strange quarks. In this review, we utilize the three-flavor NJL model to describe quark matter inside strange stars. The model exhibits spontaneous breaking of chiral symmetry, which is essential for understanding the large nucleon mass and the dynamic generation of fermion masses. Additionally, the model is noteworthy for its solvability, allowing for simple analytical results obtained in certain limiting cases.
@@ 4.3 Quasi-particle model
In view of quark confinement, the energy density in the vacuum is lower than that of free quarks. Consequently, the vacuum pressure at zero chemical potential ( ) is negative. Using the MIT bag model for reference, we can set . In this framework, quarks are treated as Fermi gas at high chemical potentials, which to some extent reflects the asymptotic freedom property of QCD. It also matches well with the results of Lattice QCD (Peshier et al., 2000, 2002; Rebhan and Romatschke, 2003).
4.4 Tolman-Oppenheimer-Volkoff equation
Due to the extremely high density of strange stars accompanied by strong gravity, the effects of spacetime curvature cannot be ignored. Consequently, the structure of strange stars has to be studied in the context of General Relativity. The Tolman-Oppenheimer-Volkoff (TOV) equation (Tolman, 1939; Oppenheimer and Volkoff, 1939), should be employed to infer the structure of such compact objects, which assumes that the interior of the star is composed of spherically symmetric ideal fluid. Using spherical coordinates of , the generalized Schwarzschild metric is (Oppenheimer and Volkoff, 1939)
where and are functions of radial coordinates. Comparing it with the spacetime interval, , the nonzero covariant components of the metric tensor can be obtained as
The energy-momentum tensor of such a ideal fluid is given by
where is the pressure and is the energy density. For a static spherically symmetric star, the pressure and energy density are functions of the r-coordinate, i.e. and . and are four-dimension velocities. Equation (7) can be rewritten as
Since the star is static, there are no spatial components for the four velocities, i.e. . Considering the normalization condition of , we have (Tolman, 1934)
The Einstein’s field equation is
where is the Ricci curvature tensor, is the scalar curvature, is the cosmological constant, is Newton’s gravitational constant and is the speed of light. For simplicity, we take hereinafter. The cosmological constant is negligible for compact stars.
The Ricci tensor is expressed as (Glendenning, 1996)
where the Christoffel symbol is defined as
Using the metric tensor of Equation (6), we can obtain the non-zero Christoffel symbols as (Glendenning, 1996)
Here the primes denote differentiation with respect to the r-coordinate. Then the nonzero components of the Ricci tensor in Equation (12) is derived as (Glendenning, 1996)
Furthermore, we have
The scalar curvature is obtained from the trace of the Ricci tensor,
The Einstein’s field equation (11) can be rewritten as,
Substituting , and into the equation, we have
Note that the last two equations are identical, hence there are only three independent equations.
The boundary conditions can be taken as the Schwarzschild metric vacuum solution, which gives (Tolman, 1939)
where is the mass of the whole star. The TOV equation can then be derived as (Oppenheimer and Volkoff, 1939)
Here the mass included inside a sphere of radius is
which can be equivalently expressed in the differential form of
Given the pressure and energy density at the center of the star, i.e. or , we can calculate the mass and radius of a compact star by solving Equations (30) and (32).
@@ 4.4 Tolman-Oppenheimer-Volkoff equation
Typical mass-radius curves of bare strange stars derived by using the MIT bag model are shown in the left panel of Figure 1 (Deb et al., 2017). On the other hand, the mass-radius curves of crusted strange stars are shown in the right panel of Figure 1 (Huang and Lu, 1997). This mass-radius relationship help us understand the maximum and minimum limits of mass, the upper limit of radius, the density distribution and even the stability condition of strange stars. We can also compare the theoretically predicted parameters with observational data to test the model and gain insights into the nature of matter under extreme conditions.
4.5 Tidal deformability and Love numbers
In the framework of General Relativity, an external tidal field perturbs the spacetime geometry around a star, leading to changes in the metric coefficients. These changes can be analytically expressed in the asymptotic region far from the star. For a spherically symmetric static star with a mass of , the metric coefficient in the local rest frame (which is asymptotically Cartesian and mass-centered coordinates) at a distance of can be given by solving the Einstein’s field equation, which is (Thorne, 1998)
where . This expression is in units of . The first term on the right hand is a general relativistic correction for the classical Newtonian gravitational potential. The symmetric and traceless tensors and represent the quadrupole moment and the external quadrupolar tidal field, respectively.
The distortion of an object caused by external gravitational forces can be described by a linear function between the quadrupole moment and the external quadrupolar tidal field (Hinderer, 2008; Flanagan and Hinderer, 2008),
where the coefficient is the so called tidal deformability. It represents the extent to which the star deforms under the influence of a specific tidal force. A commonly used dimensionless tidal deformation parameter is then defined as,
Another useful dimensionless tidal Love number () connected with the coefficient as
where is the radius of the star. is called the Love number, which represents the second spherical harmonic function (for the angular quantum number ) used in calculating the metric in Equation (33). It can be calculated as Hinderer (2008)
where is the compactness of the star. The function satisfies the differential equation of (Postnikov et al., 2010; Takátsy and Kovács, 2020)
where and are the pressure and energy density, respectively. is expressed as
where is the sound speed. and are metric functions defined in Equation (5). Taking the boundary condition as (Damour and Nagar, 2009), the tidal Love number can be conveniently calculated by solving the TOV equation.
The tidal Love number is strongly dependent on the EOS of a star. Different EOSs of neutron stars and strange quark stars result in different . Thus the tidal deformability could be a useful probe that could help distinguish between neutron stars and strange stars (Postnikov et al., 2010).
4.6 Oscillations and quasi-normal modes
Oscillations are closely relevant to the stability of stars and are sensitive to the equation of state and the composition. In this aspect, quasi-normal modes are usually discussed instead of normal modes, because the stars are practically in a system with energy losses due to gravitational radiation and other dissipative effects. The frequency of a quasi-normal mode is usually expressed as a complex number, in which the real part represents the actual oscillation frequency and the imaginary part indicates the decay rate.
Quasi-normal modes include two parts, the radial oscillation and non-radial oscillation. Radial oscillation refers to the oscillation of the compact star in the radial direction. Non-radial oscillation refers to the oscillation in a non-radial direction, which is usually described by spherical harmonics. Quasi-normal modes of non-radial oscillation are more complex and involve different mode types, such as f-mode, p-mode, and g-mode. Studying these quasi-normal modes can reveal oscillatory behaviors of compact stars under gravitational wave emissions and other dissipative effects, and help probe their internal structure.
4.6.1 radial oscillations
Radial oscillations of compact stars are investigated firstly by Chandrasekhar (1964a, b). Usually adiabatic oscillations are considered: the whole star oscillates like a retractable spring to expand and shrink periodically. Assuming a spherical symmetry and considering small-amplitude radial oscillations, we can introduce a time-dependent radial displacement of a fluid element at , which is expressed as
where represents the oscillation amplitude and is the oscillation frequency.
Under small perturbations, the contributions from nonlinear terms can be ignored. The differential equation of the radial displacement () was derived by Chandrasekhar (1964b) as
where
Here and are again the functions defined in Equation (5). is the adiabatic index determined by the equation of state,
Using the adiabatic index, can be expressed as
where is the sound speed.
To calculate , two boundary conditions must be specified. First, the fluid at the center of the star is assumed to remain at rest, i.e.
Second, the pressure perturbation vanishes at the stellar surface, which means the Lagrangian variation of the pressure should also vanish, i.e.
Under these conditions, the eigenvalue of and the corresponding radial eigenfunction of can be solved. The radial oscillation frequencies are directly linked to the mean density and elastic properties, providing information about the internal pressure and density distribution inside the star (Benvenuto and Horvath, 1991; Vaeth and Chanmugam, 1992; Jiménez and Fraga, 2019; Bora and Dev Goswami, 2021; Rather et al., 2023).
@@ 4.6 Oscillations and quasi-normal modes
In most cases, the stellar stability against radial oscillations is investigated by applying the Bardeen-Thorne-Meltzer (BTM) criterion (Bardeen et al., 1966). The BTM criterion is usually expressed as follows: when moving toward the direction of increasing central pressure along the mass-radius curve, at each extremum, one previously stable radial mode becomes unstable if the curve bends counterclockwise, while one previously unstable radial mode becomes stable if the curve bends clockwise. For bare strange quark objects which include the whole bare strange planet-bare strange star series, the conclusion on the stability is quite clear. All configurations before the mass-radius curve reaches its maximum are stable, with stars of higher central pressure unstable. It means that all the bare strange planets and bare strange dwarfs are stable.
@@
4.6.2 non-radial oscillations
Non-radial oscillations of neutron stars was initially studied by Thorne and Campolattaro (1967), which are especially important for gravitational wave emissions (Price and Thorne, 1969). For simplicity, we adopt the Cowling approximation (Cowling, 1941; McDermott et al., 1988) and ignore the gravitational perturbations in spacetime. Denoting the deviation of the fluid element from its equilibrium position as , we have (Sotani et al., 2011),
where in this subsection, and are perturbation functions of and , and is the spherical harmonics. The perturbation of the four-velocity can then be written as
Note that the energy momentum tensor satisfies (Sotani et al., 2011; Curi et al., 2022)
Assuming that the perturbations are harmonic functions of time, i.e. and , the oscillation equations in the Cowling approximation can be simplified as
To solve these equations, we again need to specify two boundary conditions. First, the Lagrangian perturbation of pressure should vanish at the stellar surface, i.e.
Second, the perturbation functions of and satisfy
at the star center (), where is a constant.
Solving this eigenvalue problem, we can determine the characteristic frequency and the corresponding characteristic function associated with the non-radial oscillations. The solutions are classified into distinct oscillation modes according to the nature of the characteristic frequencies and eigenfunctions, which mainly include:
f-modes (fundamental modes): This is the lowest order mode of non-radial oscillations. Their frequencies are typically high. They are primarily driven by the global deformation of fluid dynamics, reflecting the overall deformation of the star.
p-modes (pressure modes): For these modes, the frequency is typically high and it increases with the increasing mode order. They are primarily driven by pressure waves (sound waves) in the fluid, reflecting the pressure distribution and the speed of sound inside the star.
g-modes (gravity modes): The frequency is usually low. They are primarily driven by buoyancy forces, reflecting the density gradients and thermal gradients inside the star.
Note that under the Cowling approximation, the non-radial oscillation equations cannot effectively describe the -modes (gravitational wave modes) due to the ignoring of the metric perturbations. -modes are driven solely by gravitational waves, reflecting the spacetime undulations instead of material movements. To investigate the gravitational emissions accurately, metric perturbations in the framework of General Relativity should be considered and the coupling between the fluid dynamics equations and the Einstein’s field equations should be solved (Kokkotas and Schutz, 1992; Andersson and Kokkotas, 1996, 1998).
5 hybrid stars
A compact star is conceptually divided into five parts, the atmosphere, the outer crust, the inner crust, the outer core, and the inner core (Weber, 2005). The atmosphere is a thin layer of plasma, typically several centimeters in thickness. The outer crust is composed of atomic nuclei and free electrons, with the density being lower than g/cm3. In the inner crust, the density becomes higher but is still less than g/cm3. Atomic nucleus are disintegrated and neutrons overflow from the nucleus. In the outer core, the density increases to exceed the nuclear saturation density so that the matter is composed of neutrons, protons, electrons, and muons. In this region, all plasmas are strongly degenerate, with electrons and muons behaving like ideal Fermi gases, while protons and neutrons, among other fermionic fluids, may exist in a state of superfluidity or superconductivity. Many-body nucleon interaction models are necessary to solve for the equation of state of matter in this section under the conditions of beta equilibrium and electric neutrality. When the compact star is massive enough, it will have a special region at the center, the inner core. In this case, the density and pressure will be high enough to transform hadronic matter into quark matter (either with or without strange quarks). The transition may occur at several times of the nuclear saturation density and quark matter could coexist with hadronic matter at around the transition region. We call these stars hybrid stars. Whether a strange star or a hybrid star is more stable can be judged by comparing the energy per baryon, which is defined as the ratio of energy density to baryon number density
Currently, there is no ideal theory that can satisfactory describe the hadronic phase and the quark phase jointly. So, different models are employed to describe hadronic phase and quark phase separately. The two phases are then connected at the transition region, trying to match with each other through a particular construction, i.e. either the Maxwell construction or the Gibbs construction. Several popular models widely used to describe quark matter have been introduced in Section 4. Similarly, we have many models for hadronic matter, including the relativistic mean field (RMF) models (Walecka, 1975; Alaverdyan, 2009; Dutra et al., 2014) and the Brueckner-Hartree-Fock (BHF) approaches (Li et al., 2010; Li and Schulze, 2012; Tong et al., 2022). For example, NL3, TM1, DD-ME2 and FSU Gold models are popular RMF models, while the Akmal-Pandharipande-Ravenhall (APR) model (Akmal et al., 1998; Gusakov et al., 2005; Schneider et al., 2019), is a typical BHF approach.
Here we will focus on the transition from the hadronic phase to the quark phase. Due to unknown physics, the transition could be a smooth crossover, a sharp first-order transition involving a latent heat, or even a critical point signaling a second-order phase transition. Two methods have been engaged to math the two phases at the transition region, i.e. the Maxwell construction or the Gibbs construction.
5.1 Maxwell construction
The Maxwell construction describes a fist-order transition from hadronic matter to quark matter. In this case, the baryon number is conserved and the two phases are in equilibrium. There could be a sharp interface between the two phases so that the transition is a definite phase transitio, as shown in the left panel of Figure 2. When the pressure () is higher than the critical point (), the matter is in the quark configuration, while when is lower than , the matter is in the hadronic phase. To avoid the instabilities caused by the long-range nature of electromagnetic forces, each of the two phases have to maintain electric neutrality independently, i.e.
Under this fist-order transition, three equilibrium conditions should also be satisfied, i.e. the chemical potential equilibrium
the mechanical equilibrium
and the thermal equilibrium
where is the chemical potential of baryons, is the chemical potential of the whole system, is the chemical potential of neutron, and is temperature. The chemical potential equilibrium and mechanical equilibrium conditions are also reflected in Figure 2.
The overall EOS of a hybrid star is shown in the right panel of Figure 2. There is a “plateau” in the EOS curve, which corresponds to the transition between the hadronic phase and the quark phase. Note that the pressure is a continuous function inside the star, but there is a discontinuity in the energy density at the transition point. Such a jump in the energy density is also known as “latent heat”, which is a hallmark characteristic of a first-order phase transition.
5.2 Gibbs construction
The Gibbs construction is widely employed to describe the complex mixed matter inside compact stars (Glendenning, 1992). For the “complex” mixed phase possibly existed inside hybrid stars, keeping local electric neutrality independently is unreasonable, because the particles can interact with each other in a complicated way. Therefore, charge is conserved and electric neutrality is maintained only for the whole system, but not for each phase. This is a much weaker constraint comparing to that in the Maxwell construction.
According to the Gibbs construction, the chemical potential of each component is still equal between different phases, i.e.
where the subscript and represent the baryon phase and quark phase, respectively. is the chemical potential of electrons, which appears only in the quark phase. For the pressure and temperature , we also have
For simplicity, we could take the temperature as zero, i.e. . In this way, hadrons and quarks can not only coexist but also mix together in hybrid stars. So, the Gibbs construction is more flexible than the Maxwell construction.
The global electronic neutral condition is expressed as
where is the volume fraction of quarks that satisfies . In the mixed region, we have
Then the energy density of the mixed region is
and the baryon number density is
Comparing with the fixed parameters in the Maxwell construction, the parameters in the Gibbs construction are decided by the percentage of hadrons and quarks. In this case, the energy density is a continuous function and the conversion is smooth rather than a sharp phase boundary existed in the Maxwell construction. It is possible that there is no critical point of any phase transition, but only a crossover from hadronic matter to quark matter in the transition region. Figure 3 illustrates a typical Gibbs construction between hadronic matter and quark matter. The hadronic EOS used in the figure is from Ghosh et al. (1995) and the quark EOS is the effective mass bag model (Schertler et al., 1997, 1998). The pressure is plot as the function of two independent chemical potentials (, ). The intersection curve of the two pressure surfaces shows the solution of the Gibbs condition, where the mixed phase presents. In this transition region, all physical parameters change continuously and the free energy of the mixed phase is at a minimum.
The Gibbs construction is originally utilized to delineate the multi-phase equilibrium. The chemical potentials of each species are equal in the two phases when they coexist, which satisfies the fundamental requirement for thermodynamic equilibrium. It is not only pertinent to describe first-order phase transitions, but also can be employed to depict higher-order phase transitions and crossover phenomena under suitable circumstances.